# Searching Latent Program Spaces

### v2

@@ Abstract
General intelligence requires systems that acquire new skills efficiently and generalize beyond their training distributions. Although program synthesis approaches have strong generalization power, they face scaling issues due to large combinatorial spaces that quickly make them impractical and require human-generated DSLs or pre-trained priors to narrow this search space. On the other hand, deep learning methods have had high successes, but they lack structured test-time adaptation and rely on heavy stochastic sampling or expensive gradient updates for fine-tuning. In this work, we propose the Latent Program Network (LPN), a new architecture that builds in test-time search directly into neural models. LPN learns a latent space of implicit programs—neurally mapping inputs to outputs—through which it can search using gradients at test time. LPN combines the adaptability of symbolic approaches and the scalability of neural methods. It searches through a compact latent space at test time and bypasses the need for pre-defined domain-specific languages. On a range of programming-by-examples tasks, LPN either outperforms or matches performance compared to in-context learning and test-time training methods. Tested on the ARC-AGI benchmark, we demonstrate that LPN can both learn a compact program space and search through it at test time to adapt to novel tasks. LPN doubles its performance on out-of-distribution tasks when test-time search is switched on.
@@ 1 Introduction
The central goal of artificial intelligence has long been to create generally intelligent systems with human-like cognitive capabilities. While recent years have seen remarkable achievements in narrow AI domains, with systems achieving superhuman performance in games (Campbell et al., 2002; Silver et al., 2017; Vinyals et al., 2019) and specialized tasks (He et al., 2015; Jumper et al., 2021), we face a fundamental challenge: our systems struggle to generalize beyond their training distribution (Yu et al., 2024a) or effectively adapt to novelty (Zhang et al., 2021). This highlights a critical gap between optimizing for task-specific performance and achieving true general intelligence.
@@
In On the Measure of Intelligence, Chollet (2019) argues that traditional benchmarks that measure skills alone are insufficient for developing generally intelligent systems. Such benchmarks can be "gamed" through either unlimited prior knowledge of the task encoded by developers or massive amounts of training data, masking a system’s true ability to generalize and adapt. Instead, it is argued we need to measure skill-acquisition efficiency—how effectively a system can learn new tasks with minimal experience—and evaluate generalization capability in a way that controls for both prior knowledge and training data.
@@
By controlling for prior knowledge and experience, the Abstraction and Reasoning Corpus (ARC-AGI) (Chollet, 2019) is a benchmark that measures skill acquisition efficiency rather than pure skills. Its few-shot learning setup requires systems to adapt to novel tasks with very limited data, highlighting learning efficiency. Importantly, ARC-AGI is designed to evaluate artificial intelligence systems while calibrated to human performance. Therefore, it serves as a useful compass when comparing general intelligence between humans and machines.
@@
Current approaches to solving ARC-AGI and program synthesis benchmarks generally fall into two categories: inductive and transductive methods. (1) Inductive approaches (Parisotto et al., 2016; Devlin et al., 2017b; Butt et al., 2024) infer underlying programs from examples by generating explicit programs using a domain-specific language (DSL). However, handcrafting problem-specific DSLs is inherently unscalable for real-world applications and relies on the unrealistic assumption that a human-generated DSL is always provided. Progress on AI-generated DSLs(Ellis et al., 2021) has also been limited. In this work, we investigate methods that solve problems using only input-output example data. (2) In-context learning—also called transductive—methods (Devlin et al., 2017a; Kolev et al., 2020; Li et al., 2024a) train neural models to condition on a few examples and produce the desired output for a new input, implicitly inferring the program within the model’s activations. These methods benefit from greater scalability by removing the need for a predefined DSL. However, in-context learning struggles with consistency (Devlin et al., 2017b), i.e., being able to map the inputs to the outputs of the specification itself. Performing fine-tuning at test time is a powerful way to resolve consistency, but happens to be extremely costly on large models and is also prone to over-fitting given limited data (Li et al., 2021). Additionally, in-context learning in transformers (Vaswani et al., 2017) suffers from the quadratic complexity of self-attention (Hübotter et al., 2024), which hinders scaling specification sizes.
@@
As an attempt to get the best of both inductive and transductive worlds, we introduce the Latent Program Network (LPN), which integrates the benefits of a scalable neural architecture with program induction. By representing implicit programs in a continuous latent space, LPN allows for efficient test-time program search. LPN is more efficient than test-time fine-tuning methods like in Devlin et al. (2017a) because it performs backpropagation only through a fraction of the total parameters. Also, by performing latent program aggregation and recombination in latent space, LPN removes the quadratic cost of attention when scaling specification size.
@@ 2 Related Work
The challenge of navigating vast program spaces has driven diverse approaches to program synthesis. Early researchers focused on deductive methods, using theorem-proving techniques to construct provably correct programs based on formal specifications (Manna and Waldinger, 1980). However, the difficulty of obtaining formal specifications led to a shift toward inductive approaches (Solomonoff, 1964). These methods instead infer programs from input-output examples (Shaw et al., 1975; Summers, 1977; Biermann, 1978), making program synthesis more practical for real-world applications.
@@
However, for real-world problems, the exponential search space is a significant challenge for inductive program synthesis methods (Lee et al., 2018). Three distinct paradigms have emerged to narrow down the search space, all leveraging learning. (1) Researchers have developed differentiable programming languages that integrate symbolic reasoning with continuous optimization (Feser et al., 2016, 2017; Gaunt et al., 2017). (2) Neural networks have been trained as priors to constrain the search space and enhance the efficiency of program search techniques (Balog et al., 2016; Devlin et al., 2017b). In this paradigm of leveraging neural priors for search, large language models (LLMs) have been demonstrated to be valuable for navigating complex search spaces, including discrete program synthesis (Wang et al., 2024; Li et al., 2024d; Barke et al., 2024). CodeIt (Butt et al., 2024) leverages the pre-trained CodeT5 model (Wang et al., 2023) to guide discrete program search. (3) In-context/transductive approaches bypass intermediate program representations entirely by (meta-) training neural networks to directly map input to output examples and a new input to its corresponding output (Devlin et al., 2017a, b; Kolev et al., 2020). Neural networks offer a fully differentiable program representation, allowing program induction through gradient descent in parameter space, where network weights encode the program structure (Graves, 2014; Zaremba and Sutskever, 2014; Neelakantan et al., 2015; Kaiser and Sutskever, 2015; Kurach et al., 2015). While early research focused on learning individual programs, more recent works have broadened the scope to learning multiple programs (Devlin et al., 2017a; Kolev et al., 2020; Li et al., 2024a). Pre-trained LLMs can also be used to perform transductive reasoning, where they directly generate outputs from specifications included in the prompt—a process known as in-context learning (Gendron et al., 2023; Mitchell et al., 2023; Li et al., 2024a; Brown et al., 2020). Recent works by Hendel et al. (2023); Yang et al. (2024) demonstrate the emergence of task vectors within LLM activations. LPN advances this approach by explicitly learning and disentangling the program space from other activation information. This separation yields two benefits: enhanced program representations and more efficient program space exploration. While such neural approaches eliminate the need for formal programming languages or DSLs, they sacrifice both interpretability and the ability to perform systematic search during inference. As Devlin et al. (2017b) demonstrated, neural approaches struggle to maintain consistency with given specifications and underperform compared to neural-symbolic methods (Parisotto et al., 2016).
@@
Recent work has explored fine-tuning model parameters at test time—test-time training (TTT)—to resolve this inability to be consistent with the specification (Devlin et al., 2017a; Hottung et al., 2021b; Hübotter et al., 2024; Li et al., 2024a; Akyürek et al., 2024). TTT treats the network parameters as a program, where test-time fine-tuning becomes a search through program space via parameter optimization. Since neural networks are universal function approximators (Hornik et al., 1989), the target program is likely to exist somewhere in this parameter space. Meta-learning approaches such as MAML (Finn et al., 2017), perform gradient updates on few-shot data in the full parameter space. LEO (Rusu et al., 2018) similar to our work explore a version of MAML, introducing a bottleneck, such that the inner loop optimisation is performed in latent space. Their decoder is then a hypernetwork (Ha et al., 2016) which conditions on the latent to generate the raw parameters for inference. LPN removes the need for such a hypernetwork leveraging the in-context learning abilities of transformers to directly perform generation conditioned on the latent variable.
@@
Conditioning neural models on latent spaces has emerged as a powerful approach across diverse domains. Latent space optimization has been applied to molecule generation (Gómez-Bombarelli et al., 2018) and combinatorial optimization challenges, including the traveling salesman problem (Hottung et al., 2021a; Chalumeau et al., 2023). Recent work has extended latent-based approaches to black-box optimization through energy-based models (Yu et al., 2024b), and symbolic mathematics, where latent optimization helps balance equation complexity with accuracy (Meidani et al., 2023). In program synthesis, LEAPS (Trivedi et al., 2021) demonstrated how reinforcement learning could search latent program spaces for Karel programs (Pattis, 1994).
@@
LPN can be interpreted through the lens of semi-amortised variational inference (SVI) (Kim et al., 2018; Marino et al., 2018). SVI operates in two phases: first, it performs amortized variational inference using an encoder trained on the complete dataset. Then, for each data point, it performs additional updates to minimize the amortization gap (Gershman and Goodman, 2014)—the discrepancy between the log-likelihood and the evidence lower bound (ELBO) (Krishnan et al., 2018; Cremer et al., 2018). The LPN equivalent of these updates for test examples is performing gradient descent in the latent space.
@@ 3 Background
The objective is to generalize from the provided examples and predict the corresponding output for $x_{n+1}^{m}$. This can be done via induction or transduction (Li et al., 2024a). If we do not limit the Kolmogorov complexity of programs (Kolmogorov, 1965; Solomonoff, 1964), we can find an explanation for any given specification, whether or not it corresponds to the true program that underlies the specification. Evaluating generalization performance not only tests the ability to explain a specification but also whether it can successfully infer the underlying program in its generality and apply it to a new input. This problem bears a resemblance to few-shot learning, with the difference of having only one task. The ARC-AGI benchmark (Chollet, 2019) falls under this program synthesis formulation. Each of its tasks is composed of input-output pairs represented as 2D grid of shape up to 30x30, whose cells can take any of 10 colors.
@@ 4 Latent Program Network (LPN)
Prior work tackling programming by example with a neural approach has focused on directly training models to maximize the likelihood of decoding the correct output given a specification (in context) (Kolev et al., 2020). However, we diverge from such transduction-based methods, as they cannot inherently adapt at test time or ensure specification consistency. Instead, we explicitly factorize inference into three core components visualized in Figure 1. First, we introduce a bottleneck that encourages the network to learn an explicit representation of programs via a compact latent space. Secondly, we introduce a method for searching this latent space to explain the given data effectively. Lastly, conditioned on the latent program and a new input, we predict the output. Despite this added structure, the Latent Program Network (LPN) remains fully differentiable end-to-end. This structure provides several key benefits. Firstly, explicitly learning program representations acts as a good neural network prior for generalization. Secondly, we can use this program encoding to verify that our initial guess for a latent program explains the given data. If not, the program latent can be refined at test-time to best explain the given data. Lastly, LPN removes issues faced by transductive approaches that overfit to specification sizes (number of input-output pairs) seen during training.
@@ 4.1 Latent Program Inference
The probabilistic encoder is trained to approximate the Bayesian posterior over program latents. Specifically, it maps an input-output pair $(x,y)$ to a distribution in the latent space $q_{\phi}(z|x,y)$, representing possible programs that could explain the given input-output mapping. Using a variational approach is important because, for any given input-output pair, there exists a broad range of possible programs that map the input to the output, even when restricting to, e.g., programs of low Kolmogorov complexity (Solomonoff, 1964; Kolmogorov, 1965). We discuss LPN as semi-amortised variational inference in Appendix D. Intuitively, the encoder is trained to learn an abstract representation of programs in a continuous latent space by implicitly encoding input-output pair examples. In practice, we use a multivariate normal distribution whose mean $\mu$ and diagonal covariance $\Sigma$ parameters are inferred by the encoder. To take advantage of hardware parallelization, the encoder can process all the I/O pairs in a given specification in parallel. By encoding each pair independently and aggregating using the mean LPN is permutation invariant to the specification order, as opposed to a naive sequence model that would concatenate I/O pairs.
@@
The probabilistic decoder is responsible for mapping a latent program and an input to its expected corresponding output, directly predicting the output pixel-by-pixel instead of via a DSL. It models the distribution of possible outputs $y$ given an input $x$ and a latent $z$. Note that even if the underlying I/O mappings are deterministic, we still use a probabilistic decoding framework $p_{\theta}(y|x,z)$ to be compatible with maximum likelihood learning. Figure 18 shows the decoder generating different outputs by keeping the input fixed but varying the latent program, which in this figure represents a specific grid pattern to reproduce. In a real task, the aim of this encoder-decoder system is to learn a compressed representation of the space of possible programs we care about (e.g., in the case of ARC-AGI, this would correspond to programs that use the Core Knowledge priors (Chollet, 2019)).
@@
The encoder is trained to approximate the posterior over programs and may not encode the right abstraction given an I/O pair. Especially if the task is very novel, the encoder may fail at producing the right latent program, which, fed to the decoder, would generate the wrong output. Therefore, we include a middle stage of latent optimization where, starting from the encoder’s prediction $z$, we search for a better latent program $z^{\prime}$, one that would better explain the observed data according to the decoder $p_{\theta}$. The search process is generally denoted $z^{\prime}=f(p_{\theta},z,x,y)$ and can be implemented in several ways (c.f. section 4.2). Analogous to system 1 / system 2 thinking (Kahneman, 2011), we can think of the encoder generating an intuitive first guess as to what the observed program may be (system 1), and the latent optimization process executing a search for hypotheses that would better explain the observations (system 2). See Appendix C for test-time inference pseudo-code.

### v3

@@ Abstract
General intelligence requires systems that acquire new skills efficiently and generalize beyond their training distributions. Although program synthesis approaches have strong generalization power, they face scaling issues due to the large combinatorial spaces that quickly render them impractical, requiring human-generated DSLs or pre-trained priors to narrow this search space. On the other hand, deep learning methods have had high successes, but they lack structured test-time adaptation and rely on heavy stochastic sampling or expensive gradient updates for fine-tuning. In this work, we propose the Latent Program Network (LPN), a novel architecture that builds in test-time search directly into neural models. LPN learns a latent space of implicit programs—neurally mapping inputs to outputs—through which it can search using gradients at test time. LPN combines the adaptability of symbolic approaches and the scalability of neural methods. It searches through a compact latent space at test time and bypasses the need for pre-defined domain-specific languages. On a range of programming-by-examples tasks, LPN either outperforms or matches performance compared to in-context learning and test-time training methods. Tested on the ARC-AGI benchmark, we demonstrate that LPN can both learn a compact program space and search through it at test time to adapt to novel tasks. LPN doubles its performance on out-of-distribution tasks when test-time search is switched on.
footnotetext: Correspondence to <matthew.v.m@live.co.uk>, <clement.bonnet16@gmail.com>
@@ 1 Introduction
The central goal of artificial intelligence has long been to create generally intelligent systems with human-like cognitive capabilities. While recent years have seen remarkable achievements in narrow AI domains, with systems achieving superhuman performance in games (campbell2002deep; silver2017mastering; vinyals2019grandmaster) and specialized tasks (he2015delving; jumper2021highly), we face a fundamental challenge: our systems struggle to generalize beyond their training distribution (yu2024survey) or to effectively adapt to novelty (zhang2021understanding). This highlights a critical gap between optimizing for task-specific performance and achieving true general intelligence.
@@
In On the Measure of Intelligence, chollet2019measureintelligence argues that traditional benchmarks that measure skills alone are insufficient for developing generally intelligent systems. Such benchmarks can be "gamed" through either unlimited prior knowledge of the task encoded by developers or massive amounts of training data, masking a system’s true ability to generalize and adapt. Instead, it is argued we need to measure skill-acquisition efficiency—how effectively a system can learn new tasks with minimal experience—and evaluate generalization capability in a way that controls for both prior knowledge and training data.
@@
By controlling for prior knowledge and experience, the Abstraction and Reasoning Corpus (ARC-AGI) (chollet2019measureintelligence) is a benchmark that measures skill acquisition efficiency rather than pure skills. Its few-shot learning setup requires systems to adapt to novel tasks with very limited data, highlighting learning efficiency. Importantly, ARC-AGI is designed to evaluate artificial intelligence systems while calibrated to human performance. Therefore, it serves as a useful compass when comparing general intelligence between humans and machines.
@@
Current approaches to solving ARC-AGI and program synthesis benchmarks generally fall into two categories: inductive and transductive methods. (1) Inductive approaches (parisotto2016neuro; devlin2017robustfill; butt2024codeit) infer underlying programs from examples by generating explicit programs using a domain-specific language (DSL). However, handcrafting problem-specific DSLs is inherently unscalable for real-world applications and relies on the unrealistic assumption that a human-generated DSL is always provided. Progress on AI-generated DSLs(ellis2021dreamcoder) has also been limited. In this work, we investigate methods that solve problems using only input-output example data. (2) In-context learning—also called transductive—methods (devlin2017neural; kolev2020neural; li2024combining) train neural models to condition on a few examples and produce the desired output for a new input, implicitly inferring the program within the model’s activations. These methods benefit from greater scalability by removing the need for a predefined DSL. However, in-context learning struggles with consistency (devlin2017robustfill), i.e., being able to map the inputs to the outputs of the specification itself. Performing fine-tuning at test time is a powerful way to resolve consistency, but happens to be extremely costly on large models and is also prone to over-fitting given limited data (li2021few). Additionally, in-context learning in transformers (vaswani2017attention) suffers from the quadratic complexity of self-attention (hubotter2024efficiently), which hinders the practicality of scaling specification size.
@@
As an attempt to get the best of both inductive and transductive worlds, we introduce the Latent Program Network (LPN), which integrates the benefits of a scalable neural architecture with program induction. By representing implicit programs in a continuous latent space, LPN allows for efficient test-time program search. LPN is more efficient than test-time fine-tuning methods like in devlin2017neural because it performs backpropagation only through a fraction of the total parameters. Also, by performing latent program aggregation and recombination in latent space, LPN removes the quadratic cost of attention when scaling specification size.
@@ 2 Related Work
The challenge of navigating vast program spaces has driven diverse approaches to program synthesis. Early researchers focused on deductive methods, using theorem-proving techniques to construct provably correct programs based on formal specifications (manna1980deductive). However, the difficulty of obtaining formal specifications led to a shift toward inductive approaches (solomonoff1964formal). These methods instead infer programs from input-output examples (shaw1975inferring; summers1977methodology; biermann1978inference), making program synthesis more practical for real-world applications.
@@
However, for real-world problems, the exponential search space is a significant challenge for inductive program synthesis methods (lee2018accelerating). Three distinct paradigms have emerged to narrow down the search space, all leveraging learning. (1) Researchers have developed differentiable programming languages that integrate symbolic reasoning with continuous optimization (feser2016differentiable; feser2017neural; gaunt2017differentiable). (2) Neural networks have been trained as priors to constrain the search space and enhance the efficiency of program search techniques (balog2016deepcoder; devlin2017robustfill). In this paradigm of leveraging neural priors for search, large language models (LLMs) have been demonstrated to be valuable for navigating complex search spaces, including discrete program synthesis (wang2024planning; li2024guiding; barke2024hysynth). CodeIt (butt2024codeit) leverages the pre-trained CodeT5 model (wang2023codet5) to guide discrete program search. (3) In-context/transductive approaches bypass intermediate program representations entirely by (meta-) training neural networks to directly map input to output examples and a new input to its corresponding output (devlin2017neural; devlin2017robustfill; kolev2020neural). Neural networks offer a fully differentiable program representation, allowing program induction through gradient descent in parameter space, where network weights encode the program structure (graves2014neural; zaremba2014learning; neelakantan2015neural; kaiser2015neural; kurach2015neural). While early research focused on learning individual programs, more recent works have broadened the scope to learning multiple programs (devlin2017neural; kolev2020neural; li2024combining). Pre-trained LLMs can also be used to perform transductive reasoning, where they directly generate outputs from specifications included in the prompt—a process known as in-context learning (gendron2023large; mitchell2023comparing; li2024combining; brown2020language). Recent works by hendel2023context; yang2024task demonstrate the emergence of task vectors within LLM activations. LPN advances this approach by explicitly learning and disentangling the program space from other activation information. This separation yields two benefits: enhanced program representations and more efficient program space exploration. While such neural approaches eliminate the need for formal programming languages or DSLs, they sacrifice both interpretability and the ability to perform systematic search during inference. As devlin2017robustfill demonstrated, neural approaches struggle to maintain consistency with given specifications and underperform compared to neural-symbolic methods (parisotto2016neuro).
@@
Recent work has explored fine-tuning model parameters at test time—test-time training (TTT)—to resolve this inability to be consistent with the specification (devlin2017neural; hottung2021efficient; hubotter2024efficiently; li2024combining; akyurek2024surprising). TTT treats the network parameters as a program, where test-time fine-tuning becomes a search through program space via parameter optimization. Since neural networks are universal function approximators (hornik1989multilayer), the target program is likely to exist somewhere in this parameter space. Meta-learning approaches such as MAML (finn2017model), perform gradient updates on few-shot data in the full parameter space. LEO (rusu2018meta) similar to our work explore a version of MAML, introducing a bottleneck, such that the inner loop optimisation is performed in latent space. Their decoder is then a hypernetwork (ha2016hypernetworks) which conditions on the latent to generate the raw parameters for inference. LPN removes the need for such a hypernetwork leveraging the in-context learning abilities of transformers to directly perform generation conditioned on the latent variable.
@@
Conditioning neural models on latent spaces has emerged as a powerful approach across diverse domains. Latent space optimization has been applied to molecule generation (gomez2018automatic) and combinatorial optimization challenges, including the traveling salesman problem (hottung2021learning; chalumeau2023combinatorial). Recent work has extended latent-based approaches to black-box optimization through energy-based models (yu2024latent), and symbolic mathematics, where latent optimization helps balance equation complexity with accuracy (meidani2023snip). In program synthesis, LEAPS (trivedi2021learning) demonstrated how reinforcement learning could search latent program spaces for Karel programs (pattis1994karel).
@@
LPN can be interpreted through the lens of semi-amortised variational inference (SVI) (kim2018semi; marino2018iterative). SVI operates in two phases: first, it performs amortized variational inference using an encoder trained on the complete dataset. Then, for each data point, it performs additional updates to minimize the amortization gap (gershman2014amortized)—the discrepancy between the log-likelihood and the evidence lower bound (ELBO) (krishnan2018challenges; cremer2018inference). The LPN equivalent of these updates for test examples is performing gradient descent in the latent space.
@@ 3 Background
The objective is to generalize from the provided examples and predict the corresponding output for $x_{n+1}^{m}$. This can be done via induction or transduction (li2024combining). If we do not limit the Kolmogorov complexity of programs (kolmogorov1965three; solomonoff1964formal), we can find an explanation for any given specification, whether or not it corresponds to the true program that underlies the specification. Evaluating generalization performance not only tests the ability to explain a specification but also whether it can successfully infer the underlying program in its generality and apply it to a new input. This problem bears a resemblance to few-shot learning, with the difference of having only one task. The ARC-AGI benchmark (chollet2019measureintelligence) falls under this program synthesis formulation. Each of its tasks is composed of input-output pairs represented as 2D grid of shape up to 30x30, whose cells can take any of 10 colors.
@@ 4 Latent Program Network (LPN)
Prior work tackling programming by example with a neural approach has focused on directly training models to maximize the likelihood of decoding the correct output given a specification (in context) (kolev2020neural). However, we diverge from such transduction-based methods, as they cannot inherently adapt at test time or ensure specification consistency. Instead, we explicitly factorize inference into three core components visualized in Figure˜1. First, we introduce a bottleneck that encourages the network to learn an explicit representation of programs via a compact latent space, an architecture also used in Neural Processes (garnelo2018neural). Secondly, we introduce a method for searching this latent space to explain the given data effectively. Lastly, conditioned on the latent program and a new input, we predict the output. Despite this added structure, the Latent Program Network (LPN) remains fully differentiable end-to-end. This structure provides several key benefits. Firstly, explicitly learning program representations acts as a good neural network prior for generalization. Secondly, we can use this program encoding to verify that our initial guess for a latent program explains the given data. If not, the program latent can be refined at test-time to best explain the given data. Lastly, LPN removes issues faced by transductive approaches that overfit to specification sizes (number of input-output pairs) seen during training.
@@ 4.1 Latent Program Inference
The probabilistic encoder is trained to approximate the Bayesian posterior over program latents. Specifically, it maps an input-output pair $(x,y)$ to a distribution in the latent space $q_{\phi}(z|x,y)$, representing possible programs that could explain the given input-output mapping. Using a variational approach is important because, for any given input-output pair, there exists a broad range of possible programs that map the input to the output, even when restricting to, e.g., programs of low Kolmogorov complexity (solomonoff1964formal; kolmogorov1965three). We discuss LPN as semi-amortised variational inference in Appendix˜D. Intuitively, the encoder is trained to learn an abstract representation of programs in a continuous latent space by implicitly encoding input-output pair examples. In practice, we use a multivariate normal distribution whose mean $\mu$ and diagonal covariance $\Sigma$ parameters are inferred by the encoder. To take advantage of hardware parallelization, the encoder can process all the I/O pairs in a given specification in parallel. By encoding each pair independently and aggregating using the mean LPN is permutation invariant to the specification order, as opposed to a naive sequence model that would concatenate I/O pairs.
@@
The probabilistic decoder is responsible for mapping a latent program and an input to its expected corresponding output, directly predicting the output pixel-by-pixel instead of via a DSL. It models the distribution of possible outputs $y$ given an input $x$ and a latent $z$. Note that even if the underlying I/O mappings are deterministic, we still use a probabilistic decoding framework $p_{\theta}(y|x,z)$ to be compatible with maximum likelihood learning. Figure 20 shows the decoder generating different outputs by keeping the input fixed but varying the latent program, which in this figure represents a specific grid pattern to reproduce. In a real task, the aim of this encoder-decoder system is to learn a compressed representation of the space of possible programs we care about (e.g., in the case of ARC-AGI, this would correspond to programs that use the Core Knowledge priors (chollet2019measureintelligence)).
@@
The encoder is trained to approximate the posterior over programs and may not encode the right abstraction given an I/O pair. Especially if the task is very novel, the encoder may fail at producing the right latent program, which, fed to the decoder, would generate the wrong output. Therefore, we include a middle stage of latent optimization where, starting from the encoder’s prediction $z$, we search for a better latent program $z^{\prime}$, one that would better explain the observed data according to the decoder $p_{\theta}$. The search process is generally denoted $z^{\prime}=f(p_{\theta},z,x,y)$ and can be implemented in several ways (c.f. section 4.2). Analogous to system 1 / system 2 thinking (kahneman), we can think of the encoder generating an intuitive first guess as to what the observed program may be (system 1), and the latent optimization process executing a search for hypotheses that would better explain the observations (system 2). See Appendix˜C for test-time inference pseudo-code.